: Solve a Homogeneous DEq. ydx = 2(x + y) dy


1
Expert's answer
2021-05-07T09:40:51-0400

Given: "ydx=2(x+y)dy"

Now "ydx=2(x+y)dy\Rightarrow \frac{dx}{dy}=\frac{2(x+y)}{y}"

Observe that the given differential equation is a first-order homogeneous differential equation.

Solve the equation by using the substitution "x=vy"

"x=vy\Rightarrow \frac{dx}{dy}=v(1)+y(\frac{dv}{dy})=v+y\frac{dv}{dy}"

Then the given differential equation becomes

"\frac{dx}{dy}=\frac{2(x+y)}{y}\Rightarrow v+y\frac{dv}{dy}=\frac{2(vy+y)}{y}=\frac{2y(v+1)}{y}=2(v+1)"

"\Rightarrow y\frac{dv}{dy}=2v+2-v=v+2"

Using the separation of variables, we get

"\frac{dv}{v+2}=\frac{dy}{y}"

Integrating on both sides, we get

"\int \frac{dv}{v+2}=\int \frac{dy}{y}\Rightarrow ln(v+2)=ln(y)+ln(c)\Rightarrow ln(v+2)=ln(cy)"

"\Rightarrow v+2=cy"

Substituting "v=\frac{x}{y}" , we get

"\frac{x}{y}+2=cy\Rightarrow \frac{x+2y}{y}=cy\Rightarrow x+2y=cy^2"

Therefore, general solution to the given differential equation is

"x+2y=cy^2"